In the experiment, you determine that the exit velocity has a downward vertical component 1 m /s and the ball leaves the end of the track at height 3.12 m above floor. How much time does it take the ball to hit the floor?
Answer #1

Assuming downward velocity as negative

= yi + vyt + 1/2gt2

0 = 3.12 - 1t - 4.9t2

4.9t2 + 1t - 3.12 = 0

t = - 1 +/- 7.88 / 2 * 4.9

t = - 1 + /- 7.88 / 9.8

ignore the negative value

t = - 1 + 7.88 / 9.8

t = 0.7024 sec